[Brahms-l] Shift report 20040220 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Fri Feb 20 2004 - 07:51:53 EST
------------------------------------------------------------
Date: Fri Feb 20 07:51:52 2004, Shift: 00:00-08:00
Supervisor: DO, with: 
Log book#15 pages 62 - 62

Shift Summary:
------------------------------------------------------------
1. A promising store started around 1am, started taking 
   data but lost beams after 2mn, due to a cryo leadflow
   problem. New store at ~ 4am, still going on...

2.          MRS 40d A1050 (tr3)     FS 4d A1/2 (tr2)
    10434        139k                   157k          
    10435        218k                   190k    
    10436        166k                   150k
    10437        171k                   132k

   Changed magnet current in FS to A 1/4 
   run 10438 still running    

   pedestals: non TPC -> 10433, TPC -> 10429

3. No particular detector problems. However, the counting
   house is getting dirty and smelly (mice going after food,
   dead mice in some places, etc). Need to be fixed...
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 20 07:52:30 2004

This archive was generated by hypermail 2.1.8 : Fri Feb 20 2004 - 07:52:50 EST